OpenStudy (gabbyalicorn):

So, I was trying to find the radius of a cylinder and a question popped up in my head. IS VOLUME AND DIAMETER THE SAME THING?

OpenStudy (gabbyalicorn):

I feel like it's not but if the way its set up like this ^

OpenStudy (gabbyalicorn):

it confuses me

OpenStudy (kainui):

I see! So to understand this thing, you've probably heard for rectangles and cubes and squares that: Area = length*width Volume = length*width*height or shortened, A=lw V=lwh Looks familiar?

OpenStudy (gabbyalicorn):

yes.

OpenStudy (kainui):

Well, since A=lw and lw is part of the volume formula, we can replace it in there like this: \[V=lwh = Ah\] Kinda tricky, so now as long as we know the area of any shape with a height, we can find its volume! So a cylinder has a circle area and a height for instance. So since you know the area of a circle is \(A=\pi r^2\) you can plug this in even further: \[V = Ah = \pi r^2 h\] Now you have that equation, \[V=\pi r^2h\] you can divide both sides by \(\pi\) and \(h\) and then take the square root, and that's really where this equation comes from: \[r = \sqrt{\frac{V}{\pi h}}\]

OpenStudy (gabbyalicorn):

Oh, but that's not how to find the radius?

OpenStudy (kainui):

It isn't? The radius r is right there For a circle remember \(A=\pi r^2\) so Area = \(\pi\)* radius squared

OpenStudy (gabbyalicorn):

what about v= 1/3? My math teacher said that, that formula can only help me with cones? True?

OpenStudy (kainui):

It's similar but not the same, the volume of a cylinder is: \[V = \pi r^2 h\] The volume of a cone is: \[V = \frac{1}{3} \pi r^2h\] and the volume of a sphere is pretty similar so I'll just say it now so you don't confuse it later: \[V = \frac{4}{3} \pi r^3\] All of these things are pretty similar cause they all are sorta round in shape. It should also make some sense that you can fit a cone inside a cylinder so the volume of a cone is a third of the volume of a cylinder!

OpenStudy (gabbyalicorn):

Oh, okay. :) Thanks. I understand now.

OpenStudy (kainui):

Cool glad I could help :)